# NLP for Social Good: A Survey of Challenges, Opportunities, and Responsible Deployment

### v1

@@ Acknowledgments
This work was supported by the Novo Nordisk Foundation (grant NNF20SA0066568), which funded the research of Anders Søgaard and Antonia Karamolegkou. Antonia Karamolegkou was also supported by the Onassis Foundation - Scholarship ID: F ZP 017-2/2022-2023. Zhijing Jin was supported partly by the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039B; by the Machine Learning Cluster of Excellence, EXC number 2064/1 – Project number 390727645; and by Schmidt Sciences SAFE-AI Grant. Daryna Dementieva’s work was additionally supported by the Friedrich Schiedel TUM Fellowship.
@@
Appendix A Paper Selection: Thematic Tables of Datasets, NLP Tasks, Evaluation Metrics, and References
A.1 Health and Well-being
@@
Mental health is a key component of human health, encompassing our emotional, psychological, and social well-being SAMHSA (2023). Currently, mental health issues are a multifactorial global crisis complicated by individual risk factors and various socioeconomic and clinical factors, but NLP methods show promising potential to enhance mental healthcare Zhang et al. (2022). In this context, we define NLP tasks as activities performed by NLP techniques in roles similar to counselors and clients. In the role of counselors, NLP tools engage in several core tasks: (1) Detection and classification of mental health conditions, such as depression Giuntini et al. (2020); Mahdy et al. (2020); Khan et al. (2018) and addiction Yang et al. (2023a); Kwon et al. (2023); Ni et al. (2021), using data sources like clinical notes Panaite et al. (2022); Calvo et al. (2017) and social media posts Skaik and Inkpen (2020); Chancellor and De Choudhury (2020); Ríssola et al. (2021); (2) Responding to users by interpreting their emotional states Warikoo et al. (2022); Sabour et al. (2024); Grandi et al. (2024), generating therapeutic and empathetic responses Shen et al. (2020); Grandi et al. (2024); Sharma et al. (2023); Nazarova (2023); Zhou et al. (2025), and providing actionable feedback for support quality Min et al. (2023); Chaszczewicz et al. (2024); Althoff et al. (2016); (3) Tracking emotion and mood via time-series data analysis Čosić et al. (2024) and detecting mental health crises over time Gong et al. (2019); Yuan et al. (2023). Conversely, when NLP tools serve as clients, they typically simulate client personas from diverse backgrounds to train counselors Louie et al. (2024); Liu et al. (2025a); Hsu et al. (2025); Kampman et al. (2025). The literature for this section was selected based on the existing survey papers such as Zhang et al. (2022); Malgaroli et al. (2023) as well as keyword search (e.g., "AI for mental health", "LLM-based mental health applications", "mental health chatbots", "AI therapy").
@@
Publicly available depression-related conversations (RED, HOPE, ESC, AnnoMI-Full), expert-annotated preferences Simulating patient personas Expert evaluation on contrast with AI-like responses, linguistic authenticity, cognitive pattern authenticity, subtle emotional expression, profile adherence and personalization. Automatic evaluation on symptom severity, cognitive distortion, and overall depression Severity Liu et al. (2025a)
@@ A.7 Inequalities and Bias
LLM-Agent Interaction Logs Multi-agent bias propagation Bias-Score, fairness gap Borah and Mihalcea (2024)
BiosBias+GPT-J Causal weight-editing for debiasing Stereotype score, accuracy Cai et al. (2024)
TED-Talk / IWSLTEn–It MT Pronoun-drop attribution maps Pronoun-drop rate, TER gap Attanasio et al. (2023)
Synthetic counterfactual pairs (Iter CDA) Bias-robust text classification Bias amplification ratio, F1 Plyler and Chi (2025)
GeoWAC, Reddit, UN General Debates Region-aware bias evaluation metric Region-aware WEAT effect size, mismatchBorah et al. (2025)
CrowS-Pairs (intersectional ext.) Intrinsic bias evaluation Bias Score (Black vs White) Guo and Caliskan (2021)
@@
TV show dialogues; Cross-culture shows; LDC conversational corpora Norm Extraction; Self-Verification; Grounding AUC for grounding; human-judged best-norm selection Fung et al. (2024)
@@
A.8 Environmental Harms
Given the absence of a survey paper in this field, we began our review with research presented at the inaugural https://aclanthology.org/events/climatenlp-2024/. Additionally, we have included follow-up studies conducted by researchers within the ClimateNLP community. A significant portion of the work in this domain focuses on classifying climate-related claims, text, or stances. Detecting misinformation has emerged as a prominent topic, often tackled via external source verification and question-answering approaches. Information extraction, whether related to quantitative features or narrative insights, also plays a key role. The table 10 below highlights the key papers associated with these tasks.
@@ A.8 Environmental Harms
Appendix B Global Goals
@@
Appendix C Global Risks
@@
Appendix D Guidelines for authors

### v2

@@ Acknowledgments
This work was supported by the Novo Nordisk Foundation (grant NNF20SA0066568), which funded the research of Anders Søgaard and Antonia Karamolegkou. Antonia Karamolegkou was also supported by the Onassis Foundation - Scholarship ID: F ZP 017-2/2022-2023. Zhijing Jin was supported partly by the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039B; by the Machine Learning Cluster of Excellence, EXC number 2064/1 – Project number 390727645; and by Schmidt Sciences SAFE-AI Grant. Daryna Dementieva’s and Alexander Fraser’s work was supported by the European Research Council (ERC) under grant agreement No. 101113091 - Data4ML, an ERC Proof of Concept Grant as well as co-funded by the an ERC Advanced Grant EPICAL, grant agreement No. 101141712. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. Daryna Dementieva’s work was additionally supported by the Friedrich Schiedel TUM Think Tank Fellowship.
Appendix A Quantitative Analysis
A.1 Data Preparation
The PaperAnalyzer framework (Adauto et al., 2023) introduced a large-scale methodology for classifying research papers into socially relevant categories, drawing connections between NLP tasks, methods, and broader social goals such as the United Nations Sustainable Development Goals (SDGs). By leveraging large language models (LLMs) for annotation, PaperAnalyzer demonstrated how automated approaches can reduce the dependence on costly manual labeling while enabling scalable analysis of research trends.
Our work is inspired by this methodology and adapts it to a more focused investigation of yearly trends within the ACL Anthology. Specifically, we designed an annotation pipeline to assign papers to the set of key NLP application areas, enabling us to capture how the community’s research priorities have evolved over time.
For this study, we focused on ACL Anthology papers published between 2019 and 2024, including both main conference and workshop proceedings. This filtering yielded 47,078 papers out of the 113,207 entries in the Anthology. Across this subset, approximately 7% of papers lack abstracts; these papers were retained, with only their titles used for annotation.
A.2 Annotation of Key NLP Domains
A.2.1 Methodology
To analyze the research landscape from a social good perspective, we use a zero-shot annotation pipeline that categorized papers into nine NLP application areas: Healthcare, Education, Poverty, Peacebuilding, Environment Protection, Inclusion & Inequalities, Online Harms, Misinformation, and AI Harms. Papers could be assigned to multiple domains when appropriate, while those not fitting any category were labeled as Unrelated.
Each paper was evaluated using its title and abstract (when available). The model was explicitly instructed with the full list of domains, together with concise descriptions for each. These descriptions were initially drafted with the assistance of ChatGPT, but were subsequently reviewed and edited by the authors to ensure contextual appropriateness for the study. Our prompt can be viewed in Figure 5.
For the large-scale annotation of the full dataset, we used OpenAI’s GPT-4.1 mini (openai/gpt-4.1-mini-2025-04-14), which provided both efficiency and value at scale. All queries were executed with temperature = 0 to ensure deterministic and reproducible outputs. The pipeline was applied to the complete set of 47,078 ACL Anthology papers published between 2019 and 2024. The total cost to annotate all papers was approximately \$12 USD.
A.2.2 Results
Overall, the pipeline annotated 36,030 papers (approximately 76.5%) with at least one valid domain label, with each of these papers receiving an average of 1.82 domain assignments. During annotation, 109 papers received at least one invalid annotation. Of these, 107 papers contained a mix of valid and invalid annotations, which we handled by retaining only the valid domain labels. The remaining 2 papers contained exclusively invalid annotations and were therefore treated as Unrelated.
To examine how social good research is distributed across different publication venues, we analyzed the distribution of the key NLP domains across conference papers and workshop papers separately, as shown in Figure 4. Conferences were identified by filtering for entries where “Conference,” “Findings,” or “Annual Meeting” appeared in the booktitle field, yielding 32,886 papers. Workshop papers were identified by filtering for entries where “Workshop” appeared in the booktitle field, yielding 12,723 papers.
Figure 3: Co-occurrence heatmap of annotated NLP domains.
(a) Conference Papers
(b) Workshop Papers
Figure 4: Overview of the normalized NLP domain distributions for conference papers ($n=32,886$) and workshop papers ($n=12,723$) from 2019 to 2024.
Based on these graphs, workshops show more pronounced year-to-year fluctuations and a stronger emphasis on emerging or experimental areas. While AI Harms and Inclusion & Inequalities remain dominant in both, workshop papers show relatively higher proportions of Online Harms and Misinformation, suggesting that exploratory and early-stage research on socially sensitive topics tends to appear more frequently in workshop venues before reaching conference-level maturity.
Overall, the co-occurrence heatmap in Figure 3 confirms these broader patterns, showing that research on AI Harms and Inclusion & Inequalities often overlaps with multiple domains, making them central themes in NLP4SG. By contrast, areas such as Poverty and Peacebuilding appear more isolated, suggesting that while ethical and fairness-related work has gained strong interdisciplinary traction, other social domains remain comparatively underexplored.
LLM Annotation Prompt You will be given the title and abstract of a research paper. Your task is to annotate the paper with the relevant social impact domains. - Review the list of domains provided. - Select all domains that are relevant to the paper based on its title and abstract; you can select multiple. - If none of the domains apply, respond with "Unrelated". - Do not explain your choices. - Do not include anything other than the list of exact domain names or the word "Unrelated". **Domains**: Healthcare – Enhance physical and mental well-being by improving access, diagnosis, and support through language technologies. Education – Ensure inclusive and equitable quality education with NLP tools for learning, teaching, and accessibility. Poverty – Reduce economic hardship by enabling data-driven policies, resource distribution, and financial inclusion with NLP. Peacebuilding – Promote peaceful societies by using NLP for dynamic conflict analysis, human rights protection, and crisis response through early-warning and mediation support. Environment Protection – Address environmental risks by monitoring and analyzing them, and supporting sustainability, climate action and environmental responsibility using NLP techniques. Inclusion and Inequalities – Foster fairness and representation by using NLP to address bias, improve accessibility, and empower marginalized voices. Online Harms – Make digital spaces safe by detecting and mitigating hate speech, harassment, and harmful online behaviors. Misinformation – Safeguard information integrity by detecting, verifying, and countering false and misleading content. AI Harms – Build responsible AI by improving transparency, accountability, and fairness to prevent unintended harm from NLP systems. **Paper Title**: {paper_title} **Paper Abstract**: {paper_abstract} Please respond in the following format only: Domain; Domain; … OR Unrelated
Figure 5: Prompt used to annotate ACL Anthology papers with the key NLP research directions.
A.3 Annotation of NLP Tasks and Methods
Building on the approach proposed in Task 3 of PaperAnalyzer (Adauto et al., 2023), which analyzed NLP research for social good through its underlying tasks and methods, we aimed to reproduce a similar insight. The goal was to capture how specific techniques and problem types evolve across socially relevant research areas, complementing our domain-level annotation.
A.3.1 Methodology
We began with the set of tasks and methods previously generated from our 900-paper sample subset. We prompted Gemma 3 (google/gemma-3-27b-it) to identify and extract the tasks and methods mentioned in each paper’s title and abstract. This process yielded approximately 750 unfiltered task terms and nearly 1,000 unfiltered method terms. To consolidate these into a more structured taxonomy, we provided all extracted terms to ChatGPT and asked it to group them into canonical categories representing the most common NLP tasks and methods. We then manually reviewed each category to ensure correctness.
The resulting taxonomy included 45 distinct task categories and 49 method categories. These categories were then used to annotate the full set dataset, where we used Google’s Gemini 2.5 Flash (google/gemini-2.5-flash) to assign one or more task and method labels to each paper based on its title and abstract. The prompt used for this extraction (shown in Figure 8) instructed the model to return only relevant tasks and methods in a standardized format, ensuring consistency across the dataset. The total cost to annotate all papers was approximately \$22 USD.
A.3.2 Results
After the annotation process, only one paper contained an invalid task label, which was manually removed. Sixteen papers contained no valid method annotations; for these, the method field was replaced with the label Unknown. Overall, task coverage was nearly complete, with 46,996 papers (99.83%) successfully assigned one or more task labels and an average of 2.74 tasks per paper. Method coverage was similarly high, with 45,034 papers (95.66%) receiving at least one method label and an average of 2.22 methods per paper. Finally, to visualize the relationships between the annotated categories, we generated a Sankey diagram (Figure 6) linking the domains, tasks, and methods, illustrating how research themes and techniques intersect across the ACL Anthology corpus.
Figure 6: Diagram showing the flow of research focus from the key NLP domains to the top 15 NLP tasks and the corresponding top 10 methodological approaches.
The diagram reveals that recent NLP4SG research has increasingly centered on AI Harms and Inclusion & Inequalities, highlighting a shift toward fairness, bias mitigation, and ethical analysis of large language models. These domains are most frequently associated with Dataset Creation, Model Analysis & Interpretability and Text Classification, indicating a growing focus on building representative data, evaluating model behavior and developing practical applications for social impact.
Together, the task and method trends (Figure 7) illustrate a clear evolution in NLP4SG research. Tasks such as Model Analysis & Interpretability, Dataset Creation & Annotation, and Text Classification have gained prominence, reflecting a move toward model evaluation and analysis. Meanwhile, Transformers and Transfer Learning are gradually giving way to approaches like Prompting and In-Context Learning, highlighting the growing integration of LLM-driven, instruction-based methodologies.
(a) Tasks
(b) Methods
Figure 7: Overview of the normalized NLP task and method distributions for ACL Anthology papers from 2019 to 2024. Each plot shows trendlines for the 14 most frequent tasks and methods, selected based on their average frequency across all years.
LLM Annotation Prompt You will be given the title and abstract of a research paper. Your task is to annotate the paper with the primary NLP task(s) and method(s) that this paper is dealing with. - Review the list of tasks and methods provided. - Select all tasks and methods that are relevant to the paper based on its title and abstract; you can select multiple. - If none of the tasks or methods apply, respond with "Unknown". - Do not explain your choices. - Do not include anything other than the list of task or method labels or the word "Unknown". - Make sure to respond with the exact label for the tasks and methods. Do not modify them in any way. **Task List**: T1. Text Classification T2. Sequence Labeling T44. Web Interaction & Automation T45. Benchmarking & Comparative Analysis **Method List**: M1. Recurrent Neural Networks (RNNs, LSTMs, GRUs) M2. Transformers & Attention Mechanisms M48. Classical ML (Logistic Regression, SVM, CRF) M49. Tool Integration & API Augmentation **Paper Title**: {paper_title} **Paper Abstract**: {paper_abstract} Please respond in the following format **only**: Tasks: T#; T#; … Methods: M#; M#; …
Figure 8: Prompt used to annotate ACL Anthology papers with NLP tasks and methods.
@@ A.3.2 Results
Appendix B Paper Selection: Thematic Tables of Datasets, NLP Tasks, Evaluation Metrics, and References
B.1 Health and Well-being
@@
Mental health is a key component of human health, encompassing our emotional, psychological, and social well-being SAMHSA (2023). Currently, mental health issues are a multifactorial global crisis complicated by individual risk factors and various socioeconomic and clinical factors, but NLP methods show promising potential to enhance mental healthcare Zhang et al. (2022). In this context, we define NLP tasks as activities performed by NLP techniques in roles similar to counselors and clients. In the role of counselors, NLP tools engage in several core tasks: (1) Detection and classification of mental health conditions, such as depression Giuntini et al. (2020); Mahdy et al. (2020); Khan et al. (2018) and addiction Yang et al. (2023a); Kwon et al. (2023); Ni et al. (2021), using data sources like clinical notes Panaite et al. (2022); Calvo et al. (2017) and social media posts Skaik and Inkpen (2020); Chancellor and De Choudhury (2020); Ríssola et al. (2021); (2) Responding to users by interpreting their emotional states Warikoo et al. (2022); Sabour et al. (2024); Grandi et al. (2024), generating therapeutic and empathetic responses Shen et al. (2020); Grandi et al. (2024); Sharma et al. (2023); Nazarova (2023); Zhou et al. (2025), and providing actionable feedback for support quality Min et al. (2023); Chaszczewicz et al. (2024); Althoff et al. (2016); (3) Tracking emotion and mood via time-series data analysis Čosić et al. (2024) and detecting mental health crises over time Gong et al. (2019); Yuan et al. (2023). Conversely, when NLP tools serve as clients, they typically simulate client personas from diverse backgrounds to train counselors Louie et al. (2024); Liu et al. (2025b); Hsu et al. (2025); Kampman et al. (2025). The literature for this section was selected based on the existing survey papers such as Zhang et al. (2022); Malgaroli et al. (2023) as well as keyword search (e.g., "AI for mental health", "LLM-based mental health applications", "mental health chatbots", "AI therapy").
@@
Publicly available depression-related conversations (RED, HOPE, ESC, AnnoMI-Full), expert-annotated preferences Simulating patient personas Expert evaluation on contrast with AI-like responses, linguistic authenticity, cognitive pattern authenticity, subtle emotional expression, profile adherence and personalization. Automatic evaluation on symptom severity, cognitive distortion, and overall depression Severity Liu et al. (2025b)
@@ B.7 Inequalities and Bias
LLM‑Agent Interaction Logs Multi‑agent bias propagation Bias‑Score, fairness gap Borah and Mihalcea (2024)
BiosBias+GPT‑J Causal weight‑editing for debiasing Stereotype score, accuracy Cai et al. (2024)
TED‑Talk / IWSLTEn–It MT Pronoun‑drop attribution maps Pronoun‑drop rate, TER gap Attanasio et al. (2023)
Synthetic counterfactual pairs (Iter CDA) Bias‑robust text classification Bias amplification ratio, F1 Plyler and Chi (2025)
GeoWAC, Reddit, UN General Debates Region-aware bias evaluation metric Region‑aware WEAT effect size, mismatch% Borah et al. (2025)
CrowS‑Pairs (intersectional ext.) Intrinsic bias evaluation Bias Score (Black vs White) Guo and Caliskan (2021)
@@
TV show dialogues; Cross‑culture shows; LDC conversational corpora Norm Extraction; Self-Verification; Grounding AUC for grounding; human-judged best-norm selection Fung et al. (2024)
@@
B.8 Environmental Harms
Given the absence of a survey paper in this field, we began our review with research presented at the inaugural Workshop on Natural Language Processing Meets Climate Change (ClimateNLP 2024). Additionally, we have included follow-up studies conducted by researchers within the ClimateNLP community. A significant portion of the work in this domain focuses on classifying climate-related claims, text, or stances. Detecting misinformation has emerged as a prominent topic, often tackled via external source verification and question-answering approaches. Information extraction, whether related to quantitative features or narrative insights, also plays a key role. The table 10 below highlights the key papers associated with these tasks.
@@ B.8 Environmental Harms
Appendix C Global Goals
@@
Appendix D Global Risks
@@
Appendix E Guidelines for authors